OpenStudy (anonymous):

Determine the solution of the system. y=x+6 y=2x

OpenStudy (anonymous):

put the equations equal to each other

OpenStudy (anonymous):

That will give you the x value. Substitute back into the original problems to determine the y value

OpenStudy (anonymous):

By substitution, we obtain: 2x = x + 6 2x - x = x - x + 6 x = 6 Hence, we have (6,12) as the solution...

OpenStudy (anonymous):

6+6=>12 for y 2(6)=12

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i understand thanks